Kim traveled 2x + 50 miles on Monday and 3x – 10 miles on Tuesday.
Write a polynomial that represents the total distance traveled in two days.
Find the total distance if x = 20.

Answer by stanbon(75887)

Kim traveled 2x + 50 miles on Monday and 3x – 10 miles on Tuesday.
Write a polynomial that represents the total distance traveled in two days.
Find the total distance if x = 20.
---------------------------------------
Total distance = 2x+50 + 3x-10 = 5x+40
----
So, if x = 20, total distance = 5(20)+40 = 140
===================================================
Cheers,
Stan H.